# One Queue Is All You Need: Resolving Head-of-Line Blocking in Large Language Model Serving

### v1

@@ 9 Related Work
LLM Scheduling Optimization. Existing state-of-the-art LLM serving systems [57, 21, 49, 51] adopts continuous batching and a first-come-first-serve (FCFS) scheduling policy that suffers from head-of-line (HOL) blocking, which we address in QLM. FastServe [55] proposes preemptive scheduling with a Multi-Level Feedback Queue. Andes [27] defines Quality-of-Experience (QoE) for LLM serving as token delivery speed, and proposes a preemptive scheduler that maximizes QoE. QLM is the first queue management framework that optimizes end-to-end latency SLO attainment while improving LLM-serving throughput and device utilization by systematically orchestrating backend LSOs.
@@
LLM Serving Backend Optimization. Various LLM serving backend optimization techniques have been proposed to improve token generation throughput and memory cost while adapting to fine-tuning paradigms such StreamingLLM, Speculative Decoding, ChunkedAttention, FlashAttention and more [29, 45, 26, 61, 28, 13, 33, 1, 63, 36, 44, 64, 11, 23]. These backend LLM-serving optimizations are complementary to QLM as the LLM serving instance (see definition 2.3), and their impact on token generation throughput can be captured with profiling for the RCT Estimator (see section 6).
10 Conclusion
@@
We presented QLM, a novel queue management framework that orchestrates backend LSOs for SLO-oriented LLM serving. Evaluation using real-world LLM serving datasets on heterogeneous model types and GPU devices demonstrate that QLM improves end-to-end latency SLO attainment by 40–90% while improving serving throughput and device utilization by 20-400%. We further discuss extensions to QLM in Appendix A.3.
@@ Appendix A Appendix
A.1 Sensitivity Analysis on SLO Values
In Section 8, we configure SLO values that are derived from our production system requirements. To further explore the choice of SLO value on serving performance, we conduct a sensitivity study by varying the SLO values and showing the impact on the performance of QLM and other baselines for $W_{A}$. Figure 20 shows the impact of increasing the SLO scale (i.e., the ratio of SLO value to original SLO requirement) on SLO satisfaction across requests. As SLOs are relaxed, more baselines are able to meet all SLOs. For example, Earliest Deadline First (EDF) and SHPEHERD are the first to be able to meet all SLOs, when the SLO value is relaxed to 3× and 5× the original SLO value, respectively. This is expected as both of these mechanisms have an inherent notion of request ordering based on SLO values. When the SLO value is relaxed by 11×, vLLM is also to meet all SLOs. vLLM requires relatively high SLO relaxation because it adopts a simple first-come-first-serve (FCFS) scheduling policy without any notion of SLO-based ordering, which results in much worse HOL blocking compared to the other approaches.
Figure 20: Impact of varying SLO values on QLM.
A.2 GPU Utilization
QLM achieves higher SLO attainment and equal or higher throughput compared to other baseline approaches with the same number of GPU devices. Consequently, the utilization of these resources is expected to be higher. To demonstrate the improvement in resource utilization, we measure the average GPU memory utilization over time. GPU memory utilization is specifically chosen as LLM serving is primarily a memory-bound workload. Figure 21 shows the GPU memory utilization for $W_{A}$ (multi-model) and $W_{B}$ (single model) workloads. For $W_{A}$, we find that GPU utilization is significantly higher as QLM avoids repeated model swaps that keep the memory utilization low. For $W_{B}$, we find that GPU memory utilization is similar between QLM, EDF, and vLLM as all mechanisms use continuous batching to improve device memory utilization, and no model swapping is needed for the single-model setup. On the other hand, SHEPHERD uses dynamic batching and does not saturate GPU memory, leading to lower memory utilization and throughput.
Figure 21: GPU memory utilization comparison across LLM serving systems under multi-model and single-model setups.
A.3 Discussion
What happens if QLM is unable to meet SLOs? QLM’s plan generator may not be able to find an optimal virtual queue ordering if the request demand is high and the number of LLM serving instances (i.e., underlying compute resources) is insufficient. In such cases, we have three choices: (a) scale up the number of LLM serving instances by adding GPU devices, as we demonstrate in Figure  10 and Figure 13, (b) fall back to a heuristic such as Earliest Deadline First (EDF) and continue serving requests, and (c) performance admission control or rate limiting by dropping incoming requests to limit queue size. Option (a) can only be used when there are available resources, whereas Option (b) and Option (c) would lead to SLO violations.
Can new LSOs be added to QLM? QLM can be extended to support other LSOs that depend on the queue size and request ordering. For example, GPU partitioning techniques (such as Nvidia MIG [34]) can be added as an LSO with additional constraints on memory in the stochastic programming solver described in section 7. We leave the addition of extra LSOs to our future work.
How can QLM handle request priorities? QLM can also be used when strict priorities are assigned to requests instead of SLO values. In the strict priority model, $request_{1}$ would execute before $request_{2}$, if $priority(request_{1})$ < $priority(request_{2})$. With strict priority, the relative ordering of requests across priorities is pre-decided, however per-priority assignment still needs to be optimized to minimize request completion times. Consequently, the concepts of virtual queues, request groups, and the RCT estimator continue to remain useful. However, the underlying constraints of the stochastic programming model in the plan generator would need to be updated to support constraints unique to priorities.
Can SLOs be defined on just the waiting time? QLM addresses the problem of optimizing end-to-end latency SLO attainment. However, it can also be made to work with the modified objective of waiting time SLOs (i.e., excluding the execution time). In this case, the RCT estimator and plan generator can be simplified as only the waiting time ($W_{q}$) needs to be considered for each request group. Additionally, we can substitute the stochastic programming formulation with a linear programming solver as waiting times are largely deterministic, as shown in Insight #1 Section 2.4.

### v2

@@ 10. Related Work
LLM Scheduling and Orchestration. Existing state-of-the-art LLM serving systems (Yu et al., 2022; Kwon et al., 2023; HuggingFace, 2024) adopts continuous batching and a first-come-first-serve (FCFS) scheduling policy that suffers from head-of-line (HOL) blocking, which we address in QLM. FastServe (Wu et al., 2023b) proposes preemptive scheduling with a Multi-Level Feedback Queue. Andes (Liu et al., 2024) defines Quality-of-Experience (QoE) for LLM serving as token delivery speed, and proposes a preemptive scheduler that maximizes QoE. Llumnix (Sun et al., 2024) is an orchestration system across multiple LLM serving instances. QLM is the first queue management framework that optimizes SLO attainment while improving LLM-serving throughput and device utilization by systematically orchestrating backend LSOs.
@@
LLM Serving Backend Optimization. Various LLM serving backend optimization techniques have been proposed to improve token generation throughput and memory cost while adapting to fine-tuning paradigms such StreamingLLM, Speculative Decoding, ChunkedAttention, FlashAttention and more (Liu et al., 2023b; Sheng et al., 2023b; Lienhart, 2024; Zhang et al., 2023a; Liu et al., 2023a; Ge et al., 2024; Miao et al., 2023; Abhyankar et al., 2024; Zhong et al., 2024; Patel et al., 2023; Sheng et al., 2023a; Zhu et al., 2024; Fang et al., 2021; Leviathan et al., 2023). These backend LLM-serving optimizations are complementary to QLM as the LLM serving instance (see definition 2.3), and their impact on token generation throughput can be captured with profiling for the RCT Estimator (see section 6).
General ML Model-Serving Systems. Traditional model-serving systems provide functionalities such as scheduling, placement, batching, and autoscaling. Clipper (Crankshaw et al., 2017), TensorFlow-Serving (Olston et al., 2017), MArk (Zhang et al., 2019), InferLine (Crankshaw et al., 2020), SHEPHERD (Zhang et al., 2023b), and Clockwork (Gujarati et al., 2020) are some earlier work on serving traditional ML models like ResNet that are relatively small. INFaaS (Romero et al., 2021) and Cocktail (Gunasekaran et al., 2022) propose a model-less serving framework to automate the model selection and autoscaling to meet SLOs. However, they fail to consider the autoregressive property of LLMs. On the other hand, advanced autoscaling techniques are complementary to QLM.
11. Conclusion
@@
We presented QLM, a novel queue management framework that orchestrates backend LSOs for SLO-oriented LLM serving. Evaluation using real-world LLM serving datasets on heterogeneous model types and GPU devices demonstrate that QLM improves end-to-end latency SLO attainment by 40–90% while improving serving throughput and device utilization by 20-400%.
12. Acknowledgements
We thank the anonymous reviewers for providing their valuable feedback. We also thank Nick Hill, Lionel Villard, and Mudhakar Srivatsa for providing us with production requirements and insight, and Rohan Arora for infrastructure support. This work is supported by the National Science Foundation (NSF) under grant No. CCF 20-29049 and by the IBM-ILLINOIS Discovery Accelerator Institute (IIDAI). Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the NSF or IBM.
@@ Appendix A Appendix
A.1. Waiting Time Predictability
Symbol Description
$W$ Total waiting time for a request in the queue
$O$ Total number of output tokens in the queue
$\Theta$ Output token generation throughput (tokens/s)
$B$ Batch size
$\delta$ Decoding time per token
$\epsilon$ Preemption factor
GPU Total token memory capacity in GPU
$I_{i}$ Number of input tokens for the $i$-th request
$O_{i}$ Number of output tokens for the $i$-th request
$\mu_{O}$ Mean number of output tokens per request
$\sigma_{O}^{2}$ Variance of output tokens per request
$\mu_{I}$ Mean number of input tokens per request
$\sigma_{I}^{2}$ Variance of input tokens per request
n Size of the waiting queue
Table 4. Glossary of Symbols Used in the Statistical Derivation
Below, we present a statistical derivation for the underlying assumptions from the RWT estimator from Section 6.
The total number of output tokens in the queue would be the sum of output tokens of individual requests. The total waiting time for a request joining such a queue would be the time to process output tokens for requests ahead in the queue, which would be the number of output tokens divided by the token generation throughput.
(14) $W=\frac{O}{\Theta}$
The average token generation throughput (i.e. the number of output tokens generated per second) is simply the average batch size divided by time to generate each token.
(15) $\Theta=\frac{B}{\delta\times\epsilon}$
Due to continuous batching, the batch size is simply the number of tokens that can be kept in GPU memory.
(16) $B\approx\frac{\text{GPU}}{I_{i}+O_{i}}$
Simplifying the above equations, we derive the waiting time to be:
(17) $W=\frac{O\times\delta\times\epsilon\times E[I_{i}+O_{i}]}{GPU}$
Therefore the expected waiting time would be:
(18) $E[W]=\frac{E[O]\times\delta\times\epsilon\times E[I_{i}+O_{i}]}{GPU}$
As each request is independent of others and all requests originate from the same model, we assume the number of tokens for requests in the batch and queue to be i.i.d. (independent and identically distributed) variables.
Due to the i.i.d. condition, for a large queue, we can apply the Central Limit Theorem and approximate the total number of output tokens with a normal distribution.
(19) $O=\mathcal{N}(n\mu_{O},n\sigma_{O}^{2})$
(20) $E[O]=n\mu_{O}$
Similarly for a large batch, we can also apply the Central Limit Theorem to approximate the number tokens in a batch with a normal distribution.
(21) $I_{i}+O_{i}=\mathcal{N}(\mu_{I+O},\sigma_{I+O}^{2}/B)$
(22) $E[I_{i}+O_{i}]=\mu_{I}+\mu_{O}$
As remaining terms in the waiting time estimation are constant, we find that the expected waiting time in the queue can be estimated analytically.